# Beamfocusing Optimization for Near-Field Wideband Multi-User Communications

### v4

@@ II-C Problem Formulation
We aim to maximize the spectral efficiency of the considered near-field wideband communication system by jointly optimizing the analog and digital beamformers of the TTD-based hybrid beamforming architecture. According to (7), the signal-to-interference-plus-noise ratio (SINR) for decoding the desired signal of user $k$ at subcarrier $m$ is given by
$\gamma_{m,k}=\frac{|\mathbf{h}_{m,k}^{H}\mathbf{A}\mathbf{T}_{m}\mathbf{d}_{m,k}|^{2}}{\sum_{i\in\mathcal{K},i\neq k}|\mathbf{h}_{m,k}^{H}\mathbf{A}\mathbf{T}_{m}\mathbf{d}_{m,i}|^{2}+\sigma_{m,k}^{2}},$ (13)
where $\mathbf{d}_{m,k}\in\mathbb{C}^{N_{\mathrm{RF}}\times 1}$ is the vector at the $k$-th column of $\mathbf{D}_{m}$ and also the digital beamformer of user $k$ at subcarrier $m$. Therefore, by taking into account the OFDM cyclic prefix, the spectral efficiency of the considered system is given by
$R=\frac{1}{M+L_{\mathrm{CP}}}\sum_{m\in\mathcal{M}}\sum_{k\in\mathcal{K}}\log_{2}(1+\gamma_{m,k}).$ (14)
@@
$\max_{\mathbf{A},\mathbf{T}_{m},\mathbf{D}_{m}}\quad$ $\sum_{m\in\mathcal{M}}\sum_{k\in\mathcal{K}}\log_{2}(1+\gamma_{m,k})$ (15a)
@@
$\mathbf{T}_{m}=\mathrm{blkdiag}(e^{-j2\pi f_{m}\mathbf{t}_{1}},$
$\hskip 45.52458pt\dots,e^{-j2\pi f_{m}\mathbf{t}_{N_{\mathrm{RF}}}}),\forall m,$ (15d)
@@
where $P_{t}$ denotes the maximum transmit power available for each subcarrier [7] and set $\mathcal{A}$ collects indexes of non-zero entry in matrix $\mathbf{A}$. Problem (II-C) is generally challenging to solve due to the combined impact of the following reasons.
  • Firstly, the analog and digital beamformers are highly coupled in the non-convex sum-of-logarithms objective function and the maximum transmit power constraint.
  • Secondly, the non-zero entries of matrix $\mathbf{A}$ are subject to the non-convex unit-modulus constraints.
  • Thirdly, matrices $\mathbf{T}_{m}$ for all subcarriers are highly coupled with each other since they are implemented using the same set of time delays. Furthermore, the fixed interval constraint on the exponents of matrices $\mathbf{T}_{m}$ poses additional challenges for jointly optimizing all delays.
To the best of the authors’ knowledge, there are no general solutions to problem (II-C) in the presence of the above challenges. Therefore, in the following, we first propose an FDA approach to address this optimization problem. This approach has provable optimality but requires full CSI, determining a performance upper-bound of the TTD-based beamforming architectures. Then, an HTS approach is proposed to reduce both optimization and channel estimation overhead.
@@ III Fully-digital Approximation Approach
In this section, we develop an FDA approach, which directly maximizes the spectral efficiency, to solve problem (II-C). The main idea of the FDA approach is to optimize the hybrid beamformer sufficiently “close” to the optimal unconstrained fully-digital beamformer through the penalty method [26].
@@ III-A Problem Reformulation
In this subsection, we reformulate problem (II-C) into an FDA problem. To streamline the optimization process, we first remove the transmit power constraint in problem (II-C). Based on the fact that the optimal multi-user beamformers tend to use all available transmit power to maximize the special efficiency [27], we have the following lemma.
@@
$\max_{\mathbf{A},\mathbf{T}_{m},\mathbf{D}_{m}}\quad$ $\sum_{m\in\mathcal{M}}\sum_{k\in\mathcal{K}}\log_{2}(1+\tilde{\gamma}_{m,k})$ (16a)
@@
$\tilde{\gamma}_{m,k}=\frac{|\mathbf{h}_{m,k}^{H}\mathbf{A}\mathbf{T}_{m}\mathbf{d}_{m,k}|^{2}}{\sum_{i\in\mathcal{K},i\neq k}|\mathbf{h}_{m,k}^{H}\mathbf{A}\mathbf{T}_{m}\mathbf{d}_{m,i}|^{2}+\frac{\sigma_{m,k}^{2}}{P_{t}}\|\mathbf{A}\mathbf{T}_{m}\mathbf{D}_{m}\|_{F}^{2}}.$ (17)
Then, an optimal solution to problem (II-C) can be obtained by
@@
It is easy to verify that the solution in (18) satisfies all constraints in problem (II-C) and achieves the same objective value of problem (1) as the original solution $\tilde{\mathbf{A}}$, $\tilde{\mathbf{T}}_{m}$, and $\tilde{\mathbf{D}}_{m}$. Therefore, the solution in (18) guarantees the maximization of the objective function of problem (II-C) under the condition that the power constraint is satisfied with equality. ∎
@@
$\max_{\mathbf{A},\mathbf{T}_{m},\mathbf{D}_{m},\mathbf{W}_{m}}$ $\sum_{m\in\mathcal{M}}\sum_{k\in\mathcal{K}}\log_{2}(1+\hat{\gamma}_{m,k}(\mathbf{W}_{m}))$ (19a)
@@
$\hat{\gamma}_{m,k}(\mathbf{W}_{m})=\frac{|\mathbf{h}_{m,k}^{H}\mathbf{w}_{m,k}|^{2}}{\sum_{i\in\mathcal{K},i\neq k}|\mathbf{h}_{m,k}^{H}\mathbf{w}_{m,i}|^{2}+\frac{\sigma_{m,k}^{2}}{P_{t}}\|\mathbf{W}_{m}\|_{F}^{2}}.$ (20)
However, the new equality constraint in problem (III-A) still hinders the optimization of the hybrid beamformers. To address this issue, the penalty method can be employed, leading to the following optimization problem
$\max_{\mathbf{A},\mathbf{T}_{m},\mathbf{D}_{m},\mathbf{W}_{m}}$ $\sum_{m\in\mathcal{M}}\sum_{k\in\mathcal{K}}\log_{2}(1+\hat{\gamma}_{m,k}(\mathbf{W}_{m}))$
$-\frac{1}{\rho}\sum_{m\in\mathcal{M}}\|\mathbf{W}_{m}-\mathbf{A}\mathbf{T}_{m}\mathbf{D}_{m}\|_{F}^{2},$ (21a)
@@
where $\rho>0$ is the penalty factor. In this problem, the hybrid beamformer can be optimized to gradually approximate the fully-digital beamformer by gradually reducing the value of $\rho$. In particular, when $\rho\rightarrow 0$, the penalty term can be exactly zero, which yields a solution to problem (III-A). However, to ensure optimal performance, the penalty factor is typically initialized with a large value, ensuring that the original objective function is sufficiently maximized. Subsequently, the penalty factor is gradually decreased to guarantee the penalty term approaches zero, i.e., the equality constraint is achieved. Based on the results in [26], it can be proved that the penalty method guarantees the convergence to a stationary point of problem (III-A), which is also a stationary point of problem (1). In the following, we will provide solutions to problem (III-A) for the fully-connected and sub-connected architectures, respectively.
@@ III-B Solution for Fully-Connected Architectures
In problem (III-A), all constraints are related to the optimization variables $\mathbf{A}$ and $\mathbf{T}_{m},\forall m$. This observation motivates us to apply the block coordinate descent (BCD) method, which alternately optimizes each block with the other blocks fixed in each iteration, to solve this problem. In the following, we elaborate on the optimization of each block.
@@ III-B1 Subproblem With Respect to $\mathbf{A}$ and $\mathbf{T}_{m}$
$\min_{\mathbf{A},\mathbf{T}_{m}}\quad$ $\sum_{m\in\mathcal{M}}\left\|\mathbf{W}_{m}-\mathbf{A}\mathbf{T}_{m}\mathbf{D}_{m}\right\|_{F}^{2}$ (22a)
@@
$\min_{\mathbf{A},\mathbf{T}_{m},\mathbf{V}_{m}}$ $\sum_{m\in\mathcal{M}}\left(\left\|\mathbf{W}_{m}-\mathbf{V}_{m}\mathbf{D}_{m}\right\|_{F}^{2}+\frac{1}{\bar{\rho}}\|\mathbf{V}_{m}-\mathbf{A}\mathbf{T}_{m}\|_{F}^{2}\right),$ (23a)
@@
  • Subproblem With Respect to $\mathbf{A}:$ By fixing $\mathbf{T}_{m}$ and $\mathbf{V}_{m}$, problem (III-B1) reduces to
    $\min_{\mathbf{A}}\quad$ $\sum_{m\in\mathcal{M}}\left\|\mathbf{V}_{m}-\mathbf{A}\mathbf{T}_{m}\right\|_{F}^{2}$ (24a)
@@
    $=$ $\sum_{n=1}^{N_{\mathrm{RF}}}\sum_{l=1}^{N_{\mathrm{T}}}\sum_{m\in\mathcal{M}}\Big{(}\mathbf{a}_{n,l}^{H}\mathbf{a}_{n,l}+\mathbf{v}_{m,n,l}^{H}\mathbf{v}_{m,n,l}$
    $\hskip 36.98866pt-2\mathrm{Re}\left\{\mathbf{v}_{m,n,l}^{H}\mathbf{a}_{n,l}e^{-j2\pi f_{m}t_{n,l}}\right\}\Big{)},$ (25)
    where $\mathbf{v}_{m,n,l}=\mathbf{V}_{m}\left((l-1)\frac{N}{N_{\mathrm{T}}}+1:l\frac{N}{N_{\mathrm{T}}},n\right)$. There are two key observations for the above new objective function. Firstly, according to the unit-modulus constraint on $\mathbf{a}_{n,l}$, we have $\mathbf{a}_{n,l}^{H}\mathbf{a}_{n,l}=MN/N_{\mathrm{T}}$, which is a constant. Therefore, the optimization variable is only related to the last term. Secondly, vectors $\mathbf{a}_{n,l},\forall n,l,$ are no longer coupled with each other and can be optimized independently. Based on these observations, for each vector $\mathbf{a}_{n,l}$, its optimal solution can be obtained by solving the following optimization problem:
    $\max_{\mathbf{a}_{n,l}}\quad$ $\sum_{m\in\mathcal{M}}\mathrm{Re}\left\{\mathbf{v}_{m,n,l}^{H}\mathbf{a}_{n,l}e^{-j2\pi f_{m}t_{n,l}}\right\}$ (26a)
@@
    It can be readily obtained that the optimal solution is given by
    $\mathbf{a}_{n,l}=e^{j\angle\left(\sum_{m\in\mathcal{M}}\mathbf{v}_{m,n,l}e^{j2\pi f_{m}t_{n,l}}\right)}.$ (27)
  • Subproblem With Respect to $\mathbf{T}_{m}:$ Next, we optimize the TTD-based analog beamformer $\mathbf{T}_{m},\forall m,$ by fixing the other optimization variables. The resulting optimization is given by
    $\min_{\mathbf{T}_{m}}\quad$ $\sum_{m\in\mathcal{M}}\|\mathbf{V}_{m}-\mathbf{A}\mathbf{T}_{m}\|_{F}^{2}$ (28a)
@@
    $\max_{t_{n,l}}\quad$ $\sum_{m\in\mathcal{M}}\mathrm{Re}\left\{\mathbf{v}_{m,n,l}^{H}\mathbf{a}_{n,l}e^{-j2\pi f_{m}t_{n,l}}\right\}\triangleq h(t_{n,l})$ (29a)
@@
    This problem is a classical single-variable optimization problem within a fixed interval. One-dimensional search is an efficient method to solve this problem. Specifically, by defining a searching set $\mathcal{Q}=\{0,t_{\max}/(Q-1),2t_{\max}/(Q-1),\dots,t_{\max}\}$ with $Q$ searching points, a nearly-optimal solution of $t_{n,l}$ is given by
    $t_{n,l}=\arg\max_{t_{n,l}\in\mathcal{Q}}h(t_{n,l}).$ (30)
  • Subproblem With Respect to $\mathbf{V}_{m}$: The optimal solution of $\mathbf{V}_{m}$ can be obtained by minimizing objective function of problem (III-B1) without constraint, which is given by
@@
Given the above solution to each block, the proposed penalty-based method for solving problem (III-B1) is summarized in Algorithm 1. The convergence of this algorithm is guaranteed since each block is updated by the optimal closed-form solution in each iteration and the objective function is bounded from below.
Algorithm 1 Penalty-based algorithm for solving (III-B1).
@@
7:     until the fractional decrease of the objective value of problem (III-B1) falls below a predefined threshold
@@
9:  until the penalty value falls below a predefined threshold

### v6

@@ II-C Problem Formulation
We aim to maximize the spectral efficiency of the considered near-field wideband communication system by jointly optimizing the analog and digital beamformers of the TTD-based hybrid beamforming architecture. According to (7), the signal-to-interference-plus-noise ratio (SINR) for decoding the desired signal of user $k$ on subcarrier $m$ is given by
$\gamma_{m,k}=\frac{|\mathbf{h}_{m,k}^{H}\mathbf{A}\mathbf{T}_{m}\mathbf{d}_{m,k}|^{2}}{\sum_{i=1,i\neq k}^{K}|\mathbf{h}_{m,k}^{H}\mathbf{A}\mathbf{T}_{m}\mathbf{d}_{m,i}|^{2}+\sigma_{m,k}^{2}},$ (13)
where $\mathbf{d}_{m,k}\in\mathbb{C}^{N_{\mathrm{RF}}\times 1}$ is the vector at the $k$-th column of $\mathbf{D}_{m}$ and also the digital beamformer of user $k$ on subcarrier $m$. Therefore, by taking into account the OFDM cyclic prefix, the spectral efficiency of the considered system is given by
$R=\frac{1}{M+L_{\mathrm{CP}}}\sum_{m=1}^{M}\sum_{k=1}^{K}\log_{2}(1+\gamma_{m,k}).$ (14)
@@
$\max_{\mathbf{A},\mathbf{T}_{m},\mathbf{D}_{m}}\quad$ $\sum_{m=1}^{M}\sum_{k=1}^{K}\log_{2}(1+\gamma_{m,k})$ (15a)
@@
$\mathbf{T}_{m}=\mathrm{blkdiag}\big{(}e^{-j2\pi f_{m}\mathbf{t}_{1}},$
$\hskip 45.52458pt\dots,e^{-j2\pi f_{m}\mathbf{t}_{N_{\mathrm{RF}}}}\big{)},\forall m,$ (15d)
@@
where $P_{t}$ denotes the maximum transmit power available for each subcarrier [8] and set $\mathcal{A}$ collects indexes of non-zero entry in matrix $\mathbf{A}$. Problem (15) is generally challenging to solve due to the high coupling between different optimization variables and different subcarriers. To the best of the authors’ knowledge, there is no generic solution to the problem (15) in the presence of the above challenges. Therefore, in the following, we first propose an FDA approach to this optimization problem, which has provable optimality but incurs higher computational complexity and higher channel estimation overhead. This approach provides theoretical performance bounds for TTD-based beamforming architectures. Then, an HTS approach is proposed to reduce the complexity.
@@ III Fully-digital Approximation Approach
In this section, we develop an FDA approach, which directly maximizes the spectral efficiency, to solve problem (15). The main idea of the FDA approach is to optimize the hybrid beamformer sufficiently “close” to the optimal unconstrained fully-digital beamformer through the penalty method [31].
@@ III-A Problem Reformulation
In this subsection, we reformulate problem (15) into an FDA problem. To streamline the optimization process, we first remove the transmit power constraint in problem (15). Based on the fact that the optimal multi-user beamformers tend to use all available transmit power to maximize the special efficiency [32], we have the following lemma.
@@
$\max_{\mathbf{A},\mathbf{T}_{m},\mathbf{D}_{m}}\quad$ $\sum_{m=1}^{M}\sum_{k=1}^{K}\log_{2}(1+\tilde{\gamma}_{m,k})$ (16a)
@@
$\tilde{\gamma}_{m,k}=\frac{|\mathbf{h}_{m,k}^{H}\mathbf{A}\mathbf{T}_{m}\mathbf{d}_{m,k}|^{2}}{\sum_{i=1,i\neq k}^{K}|\mathbf{h}_{m,k}^{H}\mathbf{A}\mathbf{T}_{m}\mathbf{d}_{m,i}|^{2}+\frac{\sigma_{m,k}^{2}}{P_{t}}\|\mathbf{A}\mathbf{T}_{m}\mathbf{D}_{m}\|_{F}^{2}}.$ (17)
Then, an optimal solution to problem (15) can be obtained by
@@
It is easy to verify that the solution in (18) satisfies all constraints in problem (15) and achieves the same objective value of problem (16) as the original solution $\tilde{\mathbf{A}}$, $\tilde{\mathbf{T}}_{m}$, and $\tilde{\mathbf{D}}_{m}$. Therefore, the solution in (18) guarantees the maximization of the objective function of problem (15) under the condition that the power constraint is satisfied with equality. ∎
@@
$\max_{\mathbf{A},\mathbf{T}_{m},\mathbf{D}_{m},\mathbf{W}_{m}}$ $\sum_{m=1}^{M}\sum_{k=1}^{K}\log_{2}(1+\hat{\gamma}_{m,k}(\mathbf{W}_{m}))$ (19a)
@@
$\hat{\gamma}_{m,k}(\mathbf{W}_{m})=\frac{|\mathbf{h}_{m,k}^{H}\mathbf{w}_{m,k}|^{2}}{\sum_{i=1,i\neq k}^{K}|\mathbf{h}_{m,k}^{H}\mathbf{w}_{m,i}|^{2}+\frac{\sigma_{m,k}^{2}}{P_{t}}\|\mathbf{W}_{m}\|_{F}^{2}}.$ (20)
However, the new equality constraint in problem (19) still hinders the optimization of the hybrid beamformers. To address this issue, the penalty method can be employed, leading to the following optimization problem
$\max_{\mathbf{A},\mathbf{T}_{m},\mathbf{D}_{m},\mathbf{W}_{m}}$ $\sum_{m=1}^{M}\sum_{k=1}^{K}\log_{2}(1+\hat{\gamma}_{m,k}(\mathbf{W}_{m}))$
$-\frac{1}{\rho}\sum_{m=1}^{M}\|\mathbf{W}_{m}-\mathbf{A}\mathbf{T}_{m}\mathbf{D}_{m}\|_{F}^{2},$ (21a)
@@
where $\rho>0$ is the penalty factor. In this problem, the hybrid beamformer can be optimized to gradually approximate the fully-digital beamformer by gradually reducing the value of $\rho$. In particular, when $\rho\rightarrow 0$, the penalty term can be exactly zero, which yields a solution to problem (19). However, to ensure optimal performance, the penalty factor is typically initialized with a large value, ensuring that the original objective function is sufficiently maximized. Subsequently, the penalty factor is gradually decreased to guarantee the penalty term approaches zero, i.e., the equality constraint is achieved. Based on the results in [31], it can be proved that the penalty method guarantees the convergence to a stationary point of problem (19), which is also a stationary point of problem (16). In the following, we will provide solutions to problem (21) for the fully-connected and sub-connected architectures, respectively.
@@ III-B Solution for Fully-connected Architectures
In problem (21), all constraints are related to the optimization variables $\mathbf{A}$ and $\mathbf{T}_{m},\forall m$. This observation motivates us to apply the block coordinate descent (BCD) method, which alternately optimizes each block with the other blocks fixed in each iteration, to solve this problem. In the following, we elaborate on the optimization of each block.
@@ III-B1 Subproblem With Respect to $\mathbf{A}$ and $\mathbf{T}_{m}$
$\min_{\mathbf{A},\mathbf{T}_{m}}\quad$ $\sum_{m=1}^{M}\left\|\mathbf{W}_{m}-\mathbf{A}\mathbf{T}_{m}\mathbf{D}_{m}\right\|_{F}^{2}$ (22a)
@@
$\min_{\mathbf{A},\mathbf{T}_{m},\mathbf{V}_{m}}$ $\sum_{m=1}^{M}\left(\left\|\mathbf{W}_{m}-\mathbf{V}_{m}\mathbf{D}_{m}\right\|_{F}^{2}+\frac{1}{\bar{\rho}}\|\mathbf{V}_{m}-\mathbf{A}\mathbf{T}_{m}\|_{F}^{2}\right),$ (23a)
@@
By fixing $\mathbf{T}_{m}$ and $\mathbf{V}_{m}$, problem (23) reduces to
$\min_{\mathbf{A}}\quad$ $\sum_{m=1}^{M}\left\|\mathbf{V}_{m}-\mathbf{A}\mathbf{T}_{m}\right\|_{F}^{2}$ (24a)
@@
$=$ $\sum_{n=1}^{N_{\mathrm{RF}}}\sum_{l=1}^{N_{\mathrm{T}}}\sum_{m=1}^{M}\Big{(}\mathbf{a}_{n,l}^{H}\mathbf{a}_{n,l}+\mathbf{v}_{m,n,l}^{H}\mathbf{v}_{m,n,l}$
$\hskip 36.98866pt-2\Re\left\{\mathbf{v}_{m,n,l}^{H}\mathbf{a}_{n,l}e^{-j2\pi f_{m}t_{n,l}}\right\}\Big{)},$ (25)
where $\mathbf{v}_{m,n,l}=\mathbf{V}_{m}\big{(}(l-1)\frac{N}{N_{\mathrm{T}}}+1:l\frac{N}{N_{\mathrm{T}}},n\big{)}$. There are two key observations for the above new objective function. Firstly, according to the unit-modulus constraint on $\mathbf{a}_{n,l}$, we have $\mathbf{a}_{n,l}^{H}\mathbf{a}_{n,l}=MN/N_{\mathrm{T}}$, which is a constant. Therefore, the optimization variable is only related to the last term. Secondly, vectors $\mathbf{a}_{n,l},\forall n,l,$ are no longer coupled with each other and can be optimized independently. Based on these observations, for each vector $\mathbf{a}_{n,l}$, its optimal solution can be obtained by solving the following optimization problem:
$\max_{\mathbf{a}_{n,l}}\quad$ $\sum_{m=1}^{M}\Re\left\{\mathbf{v}_{m,n,l}^{H}\mathbf{a}_{n,l}e^{-j2\pi f_{m}t_{n,l}}\right\}$ (26a)
@@
It can be readily obtained that the optimal solution is
$\mathbf{a}_{n,l}=e^{j\angle\left(\sum_{m=1}^{M}\mathbf{v}_{m,n,l}e^{j2\pi f_{m}t_{n,l}}\right)}.$ (27)
Next, we optimize the TTD-based analog beamformer $\mathbf{T}_{m},\forall m,$ by fixing the other optimization variables. The resulting optimization is given by
$\min_{\mathbf{T}_{m}}\quad$ $\sum_{m=1}^{M}\|\mathbf{V}_{m}-\mathbf{A}\mathbf{T}_{m}\|_{F}^{2}$ (28a)
@@
$\max_{t_{n,l}}\quad$ $\sum_{m=1}^{M}\Re\left\{\mathbf{v}_{m,n,l}^{H}\mathbf{a}_{n,l}e^{-j2\pi f_{m}t_{n,l}}\right\}$ (29a)
@@
This problem is a classical single-variable optimization problem within a fixed interval, which can be effectively solved using one-dimensional search. Specifically, by defining a searching set $\mathcal{Q}=\{0,t_{\max}/(Q-1),2t_{\max}/(Q-1),\dots,t_{\max}\}$ with $Q$ samples, a nearly-optimal $t_{n,l}$ can be obtained as
$t_{n,l}=\arg\max_{t_{n,l}\in\mathcal{Q}}\Re\left\{\mathbf{v}_{m,n,l}^{H}\mathbf{a}_{n,l}e^{-j2\pi f_{m}t_{n,l}}\right\}.$ (30)
Furthermore, given $\mathbf{A}$ and $\mathbf{T}_{m}$, the optimal solution of $\mathbf{V}_{m}$ can be obtained by minimizing objective function of problem (23) without constraint, which is given by
@@
Given the above solution to each block, the proposed penalty-based method for solving (23) is summarized in Algorithm 1. The convergence of this algorithm is guaranteed since each block is updated by the optimal solution in each iteration and the objective function is bounded from below.
Algorithm 1 Penalty-based algorithm for solving (23).
@@
7:     until the fractional decrease of the objective value of problem (23) falls below a threshold
@@
9:  until the penalty value falls below a threshold